# Fairness-Aware Domain Generalization under Covariate and Dependence Shifts

### v1

@@ Appendix F Additional Results
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$
Methods (R, 0.93) (B, 0.85) (M, 0.81)
ERM 0.91$\pm$0.07 / 0.53$\pm$0.01 / 60.21$\pm$1.48 0.90$\pm$0.07 / 0.54$\pm$0.01 / 58.93$\pm$1.10 0.92$\pm$0.04 / 0.54$\pm$0.01 / 59.49$\pm$1.50
IRM 0.98$\pm$0.01 / 0.52$\pm$0.02 / 61.61$\pm$0.80 0.94$\pm$0.04 / 0.52$\pm$0.02 / 56.89$\pm$0.73 0.92$\pm$0.02 / 0.53$\pm$0.03 / 59.64$\pm$2.33
GDRO 0.81$\pm$0.18 / 0.56$\pm$0.02 / 58.73$\pm$2.23 0.89$\pm$0.07 / 0.55$\pm$0.03 / 59.44$\pm$1.66 0.87$\pm$0.08 / 0.55$\pm$0.02 / 62.57$\pm$0.91
Mixup 0.96$\pm$0.03 / 0.53$\pm$0.01 / 62.63$\pm$1.84 0.90$\pm$0.06 / 0.54$\pm$0.04 / 58.96$\pm$2.89 0.92$\pm$0.04 / 0.54$\pm$0.03 / 58.29$\pm$0.80
MLDG 0.96$\pm$0.03 / 0.52$\pm$0.02 / 61.81$\pm$0.53 0.90$\pm$0.08 / 0.55$\pm$0.01 / 58.11$\pm$0.13 0.93$\pm$0.02 / 0.53$\pm$0.02 / 58.27$\pm$0.47
CORAL 0.95$\pm$0.02 / 0.52$\pm$0.02 / 62.17$\pm$0.92 0.93$\pm$0.04 / 0.54$\pm$0.01 / 58.06$\pm$1.99 0.95$\pm$0.03 / 0.53$\pm$0.01 / 58.84$\pm$0.74
MMD 0.91$\pm$0.05 / 0.53$\pm$0.01 / 60.34$\pm$1.39 0.89$\pm$0.07 / 0.55$\pm$0.02 / 58.47$\pm$0.35 0.92$\pm$0.02 / 0.54$\pm$0.01 / 59.31$\pm$0.40
DANN 0.83$\pm$0.13 / 0.52$\pm$0.02 / 40.80$\pm$2.47 0.96$\pm$0.02 / 0.55$\pm$0.03 / 54.55$\pm$0.17 0.88$\pm$0.04 / 0.52$\pm$0.01 / 59.19$\pm$1.21
CDANN 0.95$\pm$0.03 / 0.52$\pm$0.01 / 57.61$\pm$0.68 0.94$\pm$0.03 / 0.54$\pm$0.02 / 56.97$\pm$1.29 0.87$\pm$0.09 / 0.52$\pm$0.02 / 59.59$\pm$1.74
DDG 0.92$\pm$0.03 / 0.52$\pm$0.01 / 56.52$\pm$0.71 0.92$\pm$0.04 / 0.54$\pm$0.04 / 58.21$\pm$1.40 0.92$\pm$0.07 / 0.53$\pm$0.02 / 60.91$\pm$2.47
MBDG 0.96$\pm$0.02 / 0.52$\pm$0.01 / 55.96$\pm$1.37 0.90$\pm$0.07 / 0.70$\pm$0.01 / 51.52$\pm$1.55 0.96$\pm$0.02 / 0.53$\pm$0.03 / 58.74$\pm$2.46
DDG-FC 0.95$\pm$0.03 / 0.52$\pm$0.01 / 54.53$\pm$1.44 0.93$\pm$0.02 / 0.53$\pm$0.03 / 59.32$\pm$0.59 0.92$\pm$0.04 / 0.52$\pm$0.01 / 60.08$\pm$1.31
MBDG-FC 0.96$\pm$0.02 / 0.55$\pm$0.02 / 55.93$\pm$1.98 0.91$\pm$0.07 / 0.54$\pm$0.03 / 55.50$\pm$0.55 0.90$\pm$0.06 / 0.53$\pm$0.02 / 57.37$\pm$2.39
EIIL 0.95$\pm$0.02 / 0.52$\pm$0.02 / 58.28$\pm$3.23 0.92$\pm$0.03 / 0.54$\pm$0.02 / 56.76$\pm$3.87 0.83$\pm$0.11 / 0.54$\pm$0.02 / 59.47$\pm$1.69
FarconVAE 0.90$\pm$0.07 / 0.53$\pm$0.03 / 60.52$\pm$0.14 0.89$\pm$0.05 / 0.55$\pm$0.04 / 60.30$\pm$0.64 0.82$\pm$0.07 / 0.56$\pm$0.01 / 60.31$\pm$0.40
FATDM 0.93$\pm$0.05 / 0.52$\pm$0.01 / 59.32$\pm$1.00 0.86$\pm$0.05 / 0.58$\pm$0.02 / 59.01$\pm$0.32 0.85$\pm$0.08 / 0.53$\pm$0.02 / 60.45$\pm$0.87
FDDG 0.99$\pm$0.00 / 0.50$\pm$0.00 / 62.01$\pm$1.87 0.96$\pm$0.01 / 0.52$\pm$0.02 / 58.37$\pm$0.67 0.92$\pm$0.02 / 0.52$\pm$0.02 / 59.49$\pm$1.93
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$
Methods (Q, 0.59) (S, 0.62) Avg
ERM 0.88$\pm$0.06 / 0.57$\pm$0.02 / 62.48$\pm$0.64 0.86$\pm$0.12 / 0.61$\pm$0.03 / 54.54$\pm$0.68 0.90 / 0.56 / 59.13
IRM 0.87$\pm$0.06 / 0.54$\pm$0.01 / 55.81$\pm$1.74 0.89$\pm$0.07 / 0.54$\pm$0.03 / 57.00$\pm$2.01 0.92 / 0.53 / 58.19
GDRO 0.86$\pm$0.05 / 0.57$\pm$0.01 / 62.92$\pm$1.17 0.77$\pm$0.08 / 0.64$\pm$0.04 / 60.44$\pm$2.86 0.84 / 0.57 / 60.82
Mixup 0.93$\pm$0.04 / 0.53$\pm$0.01 / 61.34$\pm$1.60 0.84$\pm$0.08 / 0.61$\pm$0.02 / 53.07$\pm$3.13 0.91 / 0.55 / 58.86
MLDG 0.89$\pm$0.08 / 0.56$\pm$0.02 / 62.85$\pm$2.38 0.85$\pm$0.05 / 0.59$\pm$0.03 / 54.42$\pm$0.02 0.91 / 0.55 / 59.10
CORAL 0.95$\pm$0.03 / 0.53$\pm$0.02 / 61.45$\pm$0.28 0.88$\pm$0.08 / 0.54$\pm$0.03 / 52.08$\pm$1.06 0.93 / 0.53 / 58.52
MMD 0.88$\pm$0.03 / 0.56$\pm$0.01 / 62.48$\pm$1.31 0.81$\pm$0.17 / 0.61$\pm$0.02 / 57.73$\pm$1.54 0.88 / 0.56 / 59.67
DANN 0.96$\pm$0.02 / 0.53$\pm$0.02 / 63.60$\pm$0.34 0.86$\pm$0.05 / 0.56$\pm$0.03 / 58.96$\pm$0.98 0.90 / 0.54 / 55.42
CDANN 0.97$\pm$0.02 / 0.54$\pm$0.03 / 64.25$\pm$1.25 0.74$\pm$0.16 / 0.60$\pm$0.01 / 57.73$\pm$1.89 0.89 / 0.54 / 59.23
DDG 0.89$\pm$0.07 / 0.55$\pm$0.01 / 56.68$\pm$0.87 0.84$\pm$0.07 / 0.58$\pm$0.03 / 54.91$\pm$1.33 0.90 / 0.54 / 57.44
MBDG 0.96$\pm$0.03 / 0.52$\pm$0.01 / 60.73$\pm$1.56 0.90$\pm$0.04 / 0.52$\pm$0.02 / 52.45$\pm$1.98 0.93 / 0.56 / 55.88
DDG-FC 0.92$\pm$0.02 / 0.54$\pm$0.02 / 59.90$\pm$1.75 0.90$\pm$0.05 / 0.57$\pm$0.02 / 57.45$\pm$0.08 0.92 / 0.53 / 58.26
MBDG-FC 0.94$\pm$0.04 / 0.52$\pm$0.01 / 61.04$\pm$2.31 0.91$\pm$0.06 / 0.53$\pm$0.03 / 52.57$\pm$0.92 0.92 / 0.53 / 56.48
EIIL 0.84$\pm$0.12 / 0.55$\pm$0.02 / 52.18$\pm$0.26 0.95$\pm$0.03 / 0.59$\pm$0.02 / 55.74$\pm$0.12 0.90 / 0.54 / 56.49
FarconVAE 0.97$\pm$0.02 / 0.56$\pm$0.03 / 61.30$\pm$1.14 0.86$\pm$0.10 / 0.58$\pm$0.02 / 60.70$\pm$1.48 0.89 / 0.56 / 60.62
FATDM 0.85$\pm$0.05 / 0.52$\pm$0.01 / 60.35$\pm$0.44 0.88$\pm$0.03 / 0.52$\pm$0.01 / 59.22$\pm$0.09 0.87 / 0.53 / 59.67
FDDG 0.99$\pm$0.01 / 0.50$\pm$0.00 / 59.11$\pm$0.94 0.98$\pm$0.02 / 0.53$\pm$0.01 / 60.77$\pm$0.23 0.97 / 0.51 / 59.95
@@
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$
@@
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$
@@
FDDG w/o sf 0.68$\pm$0.18 / 0.57$\pm$0.02 / 93.07$\pm$0.68 0.43$\pm$0.20 / 0.60$\pm$0.03 / 95.55$\pm$2.09 0.37$\pm$0.09 / 0.59$\pm$0.03 / 92.26$\pm$0.37
FDDG w/o $T$ 0.83$\pm$0.08 / 0.56$\pm$0.01 / 92.81$\pm$0.81 0.50$\pm$0.22 / 0.56$\pm$0.01 / 95.12$\pm$0.73 0.42$\pm$0.17 / 0.59$\pm$0.02 / 92.34$\pm$0.14
FDDG w/o fc 0.59$\pm$0.16 / 0.58$\pm$0.01 / 92.92$\pm$1.35 0.36$\pm$0.08 / 0.62$\pm$0.03 / 95.55$\pm$1.84 0.42$\pm$0.20 / 0.62$\pm$0.02 / 93.35$\pm$0.83
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$
@@
FDDG w/o sf 0.49$\pm$0.13 / 0.62$\pm$0.03 / 92.61$\pm$2.32 0.69$\pm$0.22 / 0.56$\pm$0.01 / 93.28$\pm$2.31 0.35$\pm$0.26 / 0.58$\pm$0.01 / 92.18$\pm$0.46
FDDG w/o $T$ 0.39$\pm$0.07 / 0.68$\pm$0.01 / 91.46$\pm$2.05 0.92$\pm$0.06 / 0.56$\pm$0.01 / 87.87$\pm$1.25 0.52$\pm$0.23 / 0.59$\pm$0.01 / 90.78$\pm$0.31
FDDG w/o fc 0.38$\pm$0.15 / 0.72$\pm$0.03 / 92.27$\pm$0.02 0.42$\pm$0.16 / 0.67$\pm$0.03 / 92.17$\pm$0.99 0.34$\pm$0.08 / 0.72$\pm$0.03 / 91.88$\pm$0.67
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$
@@
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$
@@
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$
Methods (R, 0.93) (B, 0.85) (M, 0.81)
FDDG w/o sf 0.95$\pm$0.02 / 0.52$\pm$0.01 / 55.78$\pm$1.01 0.97$\pm$0.01 / 0.51$\pm$0.01 / 55.30$\pm$1.08 0.95$\pm$0.03 / 0.53$\pm$0.01 / 58.29$\pm$0.80
FDDG w/o $T$ 0.95$\pm$0.03 / 0.52$\pm$0.01 / 61.36$\pm$0.42 0.91$\pm$0.06 / 0.54$\pm$0.01 / 57.67$\pm$0.82 0.89$\pm$0.05 / 0.55$\pm$0.01 / 60.68$\pm$0.31
FDDG w/o fc 0.95$\pm$0.02 / 0.52$\pm$0.02 / 63.72$\pm$0.37 0.87$\pm$0.09 / 0.55$\pm$0.01 / 58.86$\pm$0.68 0.89$\pm$0.08 / 0.54$\pm$0.01 / 60.61$\pm$0.59
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$
Methods (Q, 0.59) (S, 0.62) Avg
FDDG w/o sf 0.92$\pm$0.06 / 0.54$\pm$0.02 / 57.61$\pm$1.30 0.90$\pm$0.02 / 0.59$\pm$0.02 / 52.82$\pm$1.20 0.94 / 0.53 / 55.96
FDDG w/o $T$ 0.97$\pm$0.02 / 0.52$\pm$0.01 / 59.33$\pm$0.17 0.87$\pm$0.11 / 0.57$\pm$0.01 / 55.40$\pm$0.73 0.92 / 0.54 / 58.89
FDDG w/o fc 0.83$\pm$0.08 / 0.57$\pm$0.01 / 64.17$\pm$0.35 0.89$\pm$0.06 / 0.58$\pm$0.02 / 56.51$\pm$0.84 0.89 / 0.55 / 60.77
Trade-off between fairness-accuracy. In our algorithm, because $\lambda_{2}$ is the parameter that regularizes the fair loss, we conduct additional experiments to show the change between accuracy and fairness. Our results show that the larger (smaller) $\lambda_{2}$, the better (worse) model fairness for each domain as well as in average, but it gives worse (better) model utility. Moreover, we show additional experiment results based on choosing different $\gamma_{1}$ and $\gamma_{2}$. We observe that (1) by only increasing $\gamma_{2}$, the model towards giving unfair outcomes but higher accuracy; (2) by only increasing $\gamma_{1}$, performance on both model fairness and accuracy decreases. This may be due to the failure of disentanglement of factors. Evaluation on all datasets of fairness-accuracy trade-offs is given in Tab. 14. Results in the table are average performance over target domains.
Table 14: Trade-off between fairness-accuracy.
DP $\uparrow$ / AUC $\downarrow$ / Accuracy $\uparrow$
@@
Appendix G Limitations
In Secs. 5 and F, we empirically demonstrate the effectiveness of the proposed FDDG, wherein our method is developed based on assumptions. We assume (1) data instances can be encoded into three latent factors, (2) such factors are independent of each other, and (3) each domain shares the same content space. FDDG may not work well when data are generated with more than three factors, and each is correlated to the other. To address such limitations, studies on causal learning could be a solution. Moreover, our model relies on domain augmentation. While the results demonstrate its effectiveness, it might not perform optimally when content spaces do not completely overlap across domains. In such scenarios, a preferable approach would involve initially augmenting data by minimizing semantic gaps for each class across training domains, followed by conducting domain augmentations.
FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ Appendix E Additional Results
$DP$uparrow / $AUC_{fair}$downarrow / Accuracy $\uparrow$
Methods (R, 0.93) (B, 0.85) (M, 0.81) (Q, 0.59) (S, 0.62) Avg
ERM 0.91$\pm$0.07 / 0.53$\pm$0.01 / 60.21$\pm$1.48 0.90$\pm$0.07 / 0.54$\pm$0.01 / 58.93$\pm$1.10 0.92$\pm$0.04 / 0.54$\pm$0.01 / 59.49$\pm$1.50 0.88$\pm$0.06 / 0.57$\pm$0.02 / 62.48$\pm$0.64 0.86$\pm$0.12 / 0.61$\pm$0.03 / 54.54$\pm$0.68 0.90 / 0.56 / 59.13
IRM 0.98$\pm$0.01 / 0.52$\pm$0.02 / 61.61$\pm$0.80 0.94$\pm$0.04 / 0.52$\pm$0.02 / 56.89$\pm$0.73 0.92$\pm$0.02 / 0.53$\pm$0.03 / 59.64$\pm$2.33 0.87$\pm$0.06 / 0.54$\pm$0.01 / 55.81$\pm$1.74 0.89$\pm$0.07 / 0.54$\pm$0.03 / 57.00$\pm$2.01 0.92 / 0.53 / 58.19
GDRO 0.81$\pm$0.18 / 0.56$\pm$0.02 / 58.73$\pm$2.23 0.89$\pm$0.07 / 0.55$\pm$0.03 / 59.44$\pm$1.66 0.87$\pm$0.08 / 0.55$\pm$0.02 / 62.57$\pm$0.91 0.86$\pm$0.05 / 0.57$\pm$0.01 / 62.92$\pm$1.17 0.77$\pm$0.08 / 0.64$\pm$0.04 / 60.44$\pm$2.86 0.84 / 0.57 / 60.82
Mixup 0.96$\pm$0.03 / 0.53$\pm$0.01 / 62.63$\pm$1.84 0.90$\pm$0.06 / 0.54$\pm$0.04 / 58.96$\pm$2.89 0.92$\pm$0.04 / 0.54$\pm$0.03 / 58.29$\pm$0.80 0.93$\pm$0.04 / 0.53$\pm$0.01 / 61.34$\pm$1.60 0.84$\pm$0.08 / 0.61$\pm$0.02 / 53.07$\pm$3.13 0.91 / 0.55 / 58.86
MLDG 0.96$\pm$0.03 / 0.52$\pm$0.02 / 61.81$\pm$0.53 0.90$\pm$0.08 / 0.55$\pm$0.01 / 58.11$\pm$0.13 0.93$\pm$0.02 / 0.53$\pm$0.02 / 58.27$\pm$0.47 0.89$\pm$0.08 / 0.56$\pm$0.02 / 62.85$\pm$2.38 0.85$\pm$0.05 / 0.59$\pm$0.03 / 54.42$\pm$0.02 0.91 / 0.55 / 59.10
CORAL 0.95$\pm$0.02 / 0.52$\pm$0.02 / 62.17$\pm$0.92 0.93$\pm$0.04 / 0.54$\pm$0.01 / 58.06$\pm$1.99 0.95$\pm$0.03 / 0.53$\pm$0.01 / 58.84$\pm$0.74 0.95$\pm$0.03 / 0.53$\pm$0.02 / 61.45$\pm$0.28 0.88$\pm$0.08 / 0.54$\pm$0.03 / 52.08$\pm$1.06 0.93 / 0.53 / 58.52
MMD 0.91$\pm$0.05 / 0.53$\pm$0.01 / 60.34$\pm$1.39 0.89$\pm$0.07 / 0.55$\pm$0.02 / 58.47$\pm$0.35 0.92$\pm$0.02 / 0.54$\pm$0.01 / 59.31$\pm$0.40 0.88$\pm$0.03 / 0.56$\pm$0.01 / 62.48$\pm$1.31 0.81$\pm$0.17 / 0.61$\pm$0.02 / 57.73$\pm$1.54 0.88 / 0.56 / 59.67
DANN 0.83$\pm$0.13 / 0.52$\pm$0.02 / 40.80$\pm$2.47 0.96$\pm$0.02 / 0.55$\pm$0.03 / 54.55$\pm$0.17 0.88$\pm$0.04 / 0.52$\pm$0.01 / 59.19$\pm$1.21 0.96$\pm$0.02 / 0.53$\pm$0.02 / 63.60$\pm$0.34 0.86$\pm$0.05 / 0.56$\pm$0.03 / 58.96$\pm$0.98 0.90 / 0.54 / 55.42
CDANN 0.95$\pm$0.03 / 0.52$\pm$0.01 / 57.61$\pm$0.68 0.94$\pm$0.03 / 0.54$\pm$0.02 / 56.97$\pm$1.29 0.87$\pm$0.09 / 0.52$\pm$0.02 / 59.59$\pm$1.74 0.97$\pm$0.02 / 0.54$\pm$0.03 / 64.25$\pm$1.25 0.74$\pm$0.16 / 0.60$\pm$0.01 / 57.73$\pm$1.89 0.89 / 0.54 / 59.23
DDG 0.92$\pm$0.03 / 0.52$\pm$0.01 / 56.52$\pm$0.71 0.92$\pm$0.04 / 0.54$\pm$0.04 / 58.21$\pm$1.40 0.92$\pm$0.07 / 0.53$\pm$0.02 / 60.91$\pm$2.47 0.89$\pm$0.07 / 0.55$\pm$0.01 / 56.68$\pm$0.87 0.84$\pm$0.07 / 0.58$\pm$0.03 / 54.91$\pm$1.33 0.90 / 0.54 / 57.44
MBDG 0.96$\pm$0.02 / 0.52$\pm$0.01 / 55.96$\pm$1.37 0.90$\pm$0.07 / 0.70$\pm$0.01 / 51.52$\pm$1.55 0.96$\pm$0.02 / 0.53$\pm$0.03 / 58.74$\pm$2.46 0.96$\pm$0.03 / 0.52$\pm$0.01 / 60.73$\pm$1.56 0.90$\pm$0.04 / 0.52$\pm$0.02 / 52.45$\pm$1.98 0.93 / 0.56 / 55.88
DDG-FC 0.95$\pm$0.03 / 0.52$\pm$0.01 / 54.53$\pm$1.44 0.93$\pm$0.02 / 0.53$\pm$0.03 / 59.32$\pm$0.59 0.92$\pm$0.04 / 0.52$\pm$0.01 / 60.08$\pm$1.31 0.92$\pm$0.02 / 0.54$\pm$0.02 / 59.90$\pm$1.75 0.90$\pm$0.05 / 0.57$\pm$0.02 / 57.45$\pm$0.08 0.92 / 0.53 / 58.26
MBDG-FC 0.96$\pm$0.02 / 0.55$\pm$0.02 / 55.93$\pm$1.98 0.91$\pm$0.07 / 0.54$\pm$0.03 / 55.50$\pm$0.55 0.90$\pm$0.06 / 0.53$\pm$0.02 / 57.37$\pm$2.39 0.94$\pm$0.04 / 0.52$\pm$0.01 / 61.04$\pm$2.31 0.91$\pm$0.06 / 0.53$\pm$0.03 / 52.57$\pm$0.92 0.92 / 0.53 / 56.48
EIIL 0.95$\pm$0.02 / 0.52$\pm$0.02 / 58.28$\pm$3.23 0.92$\pm$0.03 / 0.54$\pm$0.02 / 56.76$\pm$3.87 0.83$\pm$0.11 / 0.54$\pm$0.02 / 59.47$\pm$1.69 0.84$\pm$0.12 / 0.55$\pm$0.02 / 52.18$\pm$0.26 0.95$\pm$0.03 / 0.59$\pm$0.02 / 55.74$\pm$0.12 0.90 / 0.54 / 56.49
FarconVAE 0.90$\pm$0.07 / 0.53$\pm$0.03 / 60.52$\pm$0.14 0.89$\pm$0.05 / 0.55$\pm$0.04 / 60.30$\pm$0.64 0.82$\pm$0.07 / 0.56$\pm$0.01 / 60.31$\pm$0.40 0.97$\pm$0.02 / 0.56$\pm$0.03 / 61.30$\pm$1.14 0.86$\pm$0.10 / 0.58$\pm$0.02 / 60.70$\pm$1.48 0.89 / 0.56 / 60.62
FCR 0.91$\pm$0.05 / 0.55$\pm$0.02 / 56.21$\pm$1.15 0.88$\pm$0.05 / 0.56$\pm$0.02 / 55.12$\pm$1.02 0.77$\pm$0.05 / 0.55$\pm$0.01 / 58.15$\pm$0.95 0.85$\pm$0.15 / 0.59$\pm$0.02 / 59.15$\pm$1.11 0.82$\pm$0.09 / 0.55$\pm$0.02 / 51.45$\pm$1.05 0.83 / 0.56 / 56.02
FTCS 0.92$\pm$0.06 / 0.54$\pm$0.01 / 58.23$\pm$1.15 0.84$\pm$0.02 / 0.56$\pm$0.03 / 49.23$\pm$1.21 0.80$\pm$0.07 / 0.56$\pm$0.02 / 55.48$\pm$0.48 0.89$\pm$0.15 / 0.58$\pm$0.00 / 57.15$\pm$0.51 0.84$\pm$0.05 / 0.56$\pm$0.01 / 56.15$\pm$0.15 0.86 / 0.56 / 55.25
FATDM 0.93$\pm$0.05 / 0.52$\pm$0.01 / 59.32$\pm$1.00 0.86$\pm$0.05 / 0.58$\pm$0.02 / 59.01$\pm$0.32 0.85$\pm$0.08 / 0.53$\pm$0.02 / 60.45$\pm$0.87 0.85$\pm$0.05 / 0.52$\pm$0.01 / 60.35$\pm$0.44 0.88$\pm$0.03 / 0.52$\pm$0.01 / 59.22$\pm$0.09 0.87 / 0.53 / 59.67
FEDORA 0.99$\pm$0.00 / 0.50$\pm$0.00 / 62.01$\pm$1.87 0.96$\pm$0.01 / 0.52$\pm$0.02 / 58.37$\pm$0.67 0.92$\pm$0.02 / 0.52$\pm$0.02 / 59.49$\pm$1.93 0.99$\pm$0.01 / 0.50$\pm$0.00 / 59.11$\pm$0.94 0.98$\pm$0.02 / 0.53$\pm$0.01 / 60.77$\pm$0.23 0.97 / 0.51 / 59.95
@@
$DP$uparrow / $AUC_{fair}$downarrow / Accuracy $\uparrow$
@@
$DP$uparrow / $AUC_{fair}$downarrow / Accuracy $\uparrow$
@@
w/o $E_{a}$ 0.68$\pm$0.18 / 0.57$\pm$0.02 / 93.07$\pm$0.68 0.43$\pm$0.20 / 0.60$\pm$0.03 / 95.55$\pm$2.09 0.37$\pm$0.09 / 0.59$\pm$0.03 / 92.26$\pm$0.37
w/o $T$ 0.83$\pm$0.08 / 0.56$\pm$0.01 / 92.81$\pm$0.81 0.50$\pm$0.22 / 0.56$\pm$0.01 / 95.12$\pm$0.73 0.42$\pm$0.17 / 0.59$\pm$0.02 / 92.34$\pm$0.14
w/o $\mathcal{L}_{fair}$ 0.59$\pm$0.16 / 0.58$\pm$0.01 / 92.92$\pm$1.35 0.36$\pm$0.08 / 0.62$\pm$0.03 / 95.55$\pm$1.84 0.42$\pm$0.20 / 0.62$\pm$0.02 / 93.35$\pm$0.83
$DP$uparrow / $AUC_{fair}$downarrow / Accuracy $\uparrow$
@@
w/o $E_{a}$ 0.49$\pm$0.13 / 0.62$\pm$0.03 / 92.61$\pm$2.32 0.69$\pm$0.22 / 0.56$\pm$0.01 / 93.28$\pm$2.31 0.35$\pm$0.26 / 0.58$\pm$0.01 / 92.18$\pm$0.46
w/o $T$ 0.39$\pm$0.07 / 0.68$\pm$0.01 / 91.46$\pm$2.05 0.92$\pm$0.06 / 0.56$\pm$0.01 / 87.87$\pm$1.25 0.52$\pm$0.23 / 0.59$\pm$0.01 / 90.78$\pm$0.31
w/o $\mathcal{L}_{fair}$ 0.38$\pm$0.15 / 0.72$\pm$0.03 / 92.27$\pm$0.02 0.42$\pm$0.16 / 0.67$\pm$0.03 / 92.17$\pm$0.99 0.34$\pm$0.08 / 0.72$\pm$0.03 / 91.88$\pm$0.67
$DP$uparrow / $AUC_{fair}$downarrow / Accuracy $\uparrow$
@@
$DP$uparrow / $AUC_{fair}$downarrow / Accuracy $\uparrow$
@@
$DP$uparrow / $AUC_{fair}$downarrow / Accuracy $\uparrow$
Methods (R, 0.93) (B, 0.85) (M, 0.81) (Q, 0.59) (S, 0.62) Avg
w/o $E_{a}$ 0.95$\pm$0.02 / 0.52$\pm$0.01 / 55.78$\pm$1.01 0.97$\pm$0.01 / 0.51$\pm$0.01 / 55.30$\pm$1.08 0.95$\pm$0.03 / 0.53$\pm$0.01 / 58.29$\pm$0.80 0.92$\pm$0.06 / 0.54$\pm$0.02 / 57.61$\pm$1.30 0.90$\pm$0.02 / 0.59$\pm$0.02 / 52.82$\pm$1.20 0.94 / 0.53 / 55.96
w/o $T$ 0.95$\pm$0.03 / 0.52$\pm$0.01 / 61.36$\pm$0.42 0.91$\pm$0.06 / 0.54$\pm$0.01 / 57.67$\pm$0.82 0.89$\pm$0.05 / 0.55$\pm$0.01 / 60.68$\pm$0.31 0.97$\pm$0.02 / 0.52$\pm$0.01 / 59.33$\pm$0.17 0.87$\pm$0.11 / 0.57$\pm$0.01 / 55.40$\pm$0.73 0.92 / 0.54 / 58.89
w/o $\mathcal{L}_{fair}$ 0.95$\pm$0.02 / 0.52$\pm$0.02 / 63.72$\pm$0.37 0.87$\pm$0.09 / 0.55$\pm$0.01 / 58.86$\pm$0.68 0.89$\pm$0.08 / 0.54$\pm$0.01 / 60.61$\pm$0.59 0.83$\pm$0.08 / 0.57$\pm$0.01 / 64.17$\pm$0.35 0.89$\pm$0.06 / 0.58$\pm$0.02 / 56.51$\pm$0.84 0.89 / 0.55 / 60.77
Figure 8: Tracking the change of reconstruction loss using the B/W/L domains of the FairFace dataset.
Sensitive analysis on slacks $\gamma_{1}$ and $\gamma_{2}$ We show additional experiment results by choosing different $\gamma_{1}$ and $\gamma_{2}$ of Algorithm 1 in Tab. 15. We observe that (1) by only increasing $\gamma_{2}$, the model towards giving unfair outcomes but higher accuracy; (2) by only increasing $\gamma_{1}$, performance on both model fairness and accuracy decreases. This may be due to the failure of disentanglement of factors.
Table 15: Sensitiveness of slacks.
$DP$uparrow / $AUC_{fair}$downarrow / Accuracy $\uparrow$
@@
Appendix F Limitations
In Secs. 6 and E, we empirically demonstrate the effectiveness of the proposed FEDORA, wherein our method is developed based on assumptions. We assume (1) data instances can be encoded into three latent factors, (2) such factors are independent of each other, and (3) each domain shares the same semantic space. FEDORA may not work well when data are generated with more than three factors or such factors are correlated to each other. Studies on causal learning could be a solution to address such limitations. Moreover, our model relies on domain augmentation. While the results demonstrate its effectiveness, it might not perform optimally when semantic spaces do not completely overlap across domains. In such scenarios, a preferable approach would involve initially augmenting data by minimizing semantic gaps for each class across training domains, followed by conducting domain augmentations.